OpenStudy (awesome781):

A bowling ball with a mass of 7.0 kg strikes a pin that has a mass of 2.0 kg. The pin flies forward with a velocity of 6.0 m/s, and the ball continues forward at 4.0 m/s. What was the original velocity of the ball?

OpenStudy (fifciol):

 momentum must be conserved(any external forces act on the objects) \[m_1v+0=m_1v_1+m_2v_2\]

OpenStudy (awesome781):

So how do I plug the numbers?

OpenStudy (fifciol):

7*v=7*4+6*2

OpenStudy (awesome781):

I got 5.7 m/s as my answer....

OpenStudy (fifciol):

that's correct

OpenStudy (awesome781):

Ok...thanks again!!!!

OpenStudy (fifciol):

yw